# On complexity of restricted fragments of Decision DNNF

### v1

@@ 3.1 An approach to proving $\wedge_{d}$-obbd lower bounds
Let the notation and premises be as in Lemma 2. In addition, assume that $\mathcal{S}(B)|_{\bf g}\neq\emptyset$. Let $X=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\bigcup_{u\in L_{\bf g}}\mathsf{var}(B_{u})$. If $L({\bf g})\neq\emptyset$ then $\mathcal{S}(B)|_{\bf g}=X^{\{0,1\}}\times\prod_{u\in L({\bf g})}\mathcal{S}(B_{u})$. Otherwise, $\mathcal{S}(B)|_{\bf g}=X^{\{0,1\}}$.
Lemma 3 is proved by an elementary though lengthy inductive exploration of the structure of $\wedge_{d}$-obdd. A complete proof is postponed to the Appendix.
We are now ready to outline a methodology for obtaining a lower bound for a $\wedge_{d}$-obdd $B$ obeying an order $\pi$. Like in the case for obdds, we introduce a large set $\mathcal{F}$ set of fooling assignments and seek to prove that $|B|\geq|\mathcal{F}|$. Since, we are dealing with a tree $T({\bf g})$ rather than $P({\bf g})$, we make an extra effort of identifying a leaf $u({\bf g})$ of $T({\bf g})$. Having done that, we prove, like in case of obdd, that for two distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$, $u({\bf g}_{1})\neq u({\bf g}_{2})$ thus implying the desired lower bound $|B|\geq|\mathcal{F}|$.
In order to identify $u({\bf g})$, we introduce a set $UB({\bf g})\subseteq\mathsf{var}(B)\setminus\mathsf{var}({\bf g})$ which we refer to as the unbreakable set for ${\bf g}$. This is a set of at least two variables and we prove that $\mathcal{S}(B)|_{\bf g}$ does not break $UB({\bf g})$ (hence the name).
Next, we employ Lemma 3 to claim that either $UB({\bf g})\subseteq\mathsf{var}(B_{u})$ for some $u\in L({\bf g})$ or $UB({\bf g})\subseteq X$ where $X$ is as in the statement of Lemma 3. The latter possibility is ruled outr since $UB({\bf g})$ is of size at least $2$ and $\mathsf{var}(X)$ can be further broken into singletons. We thus set $u({\bf g})$ to be the $u\in L({\bf g})$ such tha $UB({\bf g})\subseteq\mathsf{var}(B_{u})$.
For the last stage of the proof, if it happens that $u({\bf g}_{1})=u({\bf g}_{2})=u$ then $UB({\bf g}_{1})\cup UB({\bf g}_{2})\subseteq\mathsf{var}(B_{u})$ simply by definition. It thus follows from Lemma 3 that $\mathcal{S}|_{{\bf g}_{1}}$ and $\mathcal{S}|_{{\bf g}_{2}}$ have the same projections to $UB({\bf g}_{1})\cup UB({\bf g}_{2})$. We prove that these projections are distinct and hence establish that $u({\bf g}_{1})\neq u({\bf g}_{2})$.
@@ 3.2 $\wedge_{d}$-obdd lower bounds and separations
In [3] we established an xp lower bound for $\wedge_{d}$-obdds representing cnfs of bounded incidence treewidth. We prove this result here using the approach outlined in Section 3.1.
@@
There is a constant $c$ and an infinite set of values $k$ such that for each $k$ there is an infinite set $\Phi$ of cnfs such that for each $\varphi\in\Phi$ $itw(\varphi)\leq k$ and $\wedge_{d}$-obdd$(\varphi)\geq|\mathsf{var}(\varphi)|^{c\cdot k}$.
Thus a natural restriction on $\wedge_{d}$-fbdd separates primal [11] and incidence treewidth. This makes $\wedge_{d}$-obdd an interesting object to study in greater detail. We establish the relationship of $\wedge_{d}$-obdd with two closely related formalisms obdd and fbdd by proving exponential separations as specified below (for both statements, the lower bounds are proved using the approach as in Section 3.1).
@@
For the corresponding quasipolynomial separation, take a class of cnfs of $ptw$ at most $k$ that require fbdd of size $n^{\Omega(k)}$ (see [15] for such a class of cnfs). Let $k=\log n$ where $n$ is the number of variables. Then the resulting class of cnfs will require $n^{\Omega(\log(n)}$ size fbdds for their representation but can be represented by a polynomial size $\wedge_{d}$-obdd according to [11].
@@
Theorem 6 looks somewhat surprising because equipping fbdds with the $\wedge$ gates leads only to quasipolynomial gains in size [1]
@@
$E(G)=\{\{u_{1},u_{2}\},\{u_{2},u_{3}\},\{u_{3},u_{4}\},\{u_{4},u_{5}\},\{u_{5},u_{1}\},\{u_{1},u_{3}\}\}$. In other words, $G$ is a cycle of $5$ vertices with an extra chord. Then $V(G^{*})=\{u_{1}[1],\dots,u_{5}[1],u_{1}[2],\dots,u_{5}[2]\}$ and $E(G^{*})=\{\{u_{1}[1],u_{2}[2]\},\{u_{2}[1],u_{3}[2]\},\{u_{3}[1],u_{4}[2]\},\{u_{4}[1],u_{5}[2]\},\{u_{5}[1],u_{1}[2]\},\{u_{1}[1],u_{3}[2]\},\\ \{u_{1}[2],u_{2}[1]\},\{u_{2}[2],u_{3}[1]\},\{u_{3}[2],u_{4}[1]\},\{u_{4}[2],u_{5}[1]\},\{u_{5}[2],u_{1}[1]\},\{u_{1}[2],u_{3}[1]\}\}$. $\varphi(G)$ and $\varphi(G^{*})$ simply turn the edges $\{u,v\}$ of the respective graphs into clauses $(u\vee v)$. $\psi(G)$ is obtained from $\varphi(G^{*})$ by adding clauses ($(\neg u_{1}[1]\vee\dots\vee\neg u_{5}[1])$ and ($(\neg u_{1}[2]\vee\dots\vee\neg u_{5}[2])$
@@
We will use cnfs with junction variables for Theorems 5 and 6. This approach is used in [20] for separation of obdds and fbdds. The rough idea is that a ’hard’ function is ’split’ into two ’easy’ ones using the junction variable. The fbdd, due to its flexibility, can adjust and represent each ’easy’ function with the suitable variable ordering. On the other hand, obdd is rigid in the sense that it needs to use the same ordering for both ’easy’ functions and the ordering leads to an exponential size representation for one of them.
In order to proceed, we need to delve deeper into the structure of matchings of graph $G^{*}$. This is done in the next definition and two statements that follow.
@@
Let $M$ be a matching of $G^{*}$ between $V^{\prime}$ and $V^{\prime\prime}$. We say that $V^{\prime},V^{\prime\prime}$ is the neat partition of $V(M)$ if there are $U,W\in V(G)$ such that,say $V^{\prime}=U[1]$ and $V^{\prime\prime}=W[2]$.
Let $\pi^{*}$ be a linear order of $V(G^{*})$. We say that $M$ neatly crosses $\pi^{*}$ if there a prefix $\pi^{*}_{0}$ of $\pi^{*}$ such that $\pi^{*}_{0}\cap V(M)$ and $(\pi^{*}\setminus\pi^{*}_{0})\cap V(M)$ form a neat partition of $M$.
In other words $M$ neatly crossing $\pi^{*}$ means that $M$ crosses $\pi^{*}$ with a witnessing prefix satisfying the constraint as in the definition. In the next statement, we demonstrate that if each linear order of $G$ is crossed by a large induced matching then every linear order of $G^{*}$ is neatly crossed by a large induced matching. The statement is formulated in terms of parameter lsim-width defined earlier.
@@
Every linear order of $G^{*}$ is neatly crossed by an induced matching of size at least $lsimw(G)$
@@
Let $u[1]$ be the last vertex of $\pi^{*}_{0}$. Let $w[2]\in W_{1}[2]$. Note that by definition of $\pi_{0}$, $u\in\pi_{0}$ while $w$ is not. It follows from $u[ind(u)]<w[ind(w)]$ according to $\pi^{*}$. Now $ind(u)=1$ by definition of $U_{1}$. That is $u[1]$ occurs before $w[ind(w)]$ and surely before $w[3-ind(w)]$ in $\pi^{*}$. As $u[1]$ is the last element of $\pi^{*}_{0}$, both occurrences of $w$ are outside of $\pi^{*}_{0}$. ∎
@@
For $i\in\{1,2\}$, Let $E^{*}_{i}=E(G[E_{i}]^{*})$. It is not hard to see that $E^{*}_{1},E^{*}_{2}$ partition $E(G^{*})$ and $G^{*}[E^{*}_{i}]=G[E_{i}]^{*}$ for each $i\in\{1,2\}$. Let $M^{*}$ be an induced matching of size at least $lsimw(G)/2$ neatly crossing $\pi^{*}$ existing by Lemma 7. For each $i\in\{1,2\}$, let $M^{*}_{i}=M^{*}\cap E^{*}_{i}$. Clearly, both $M^{*}_{1}$ and $M^{*}_{2}$ cross $\pi^{*}$ ans least one of them is of size at least half the size of $M^{*}$ that is at least $lsimw(G)/4$. ∎
@@
Note that Lemma 8 does not make any statements regarding $lsimw(G[E_{1}])$ and $lsimw(G[E_{2}])$. Indeed, if $G$ is an $n\times n$ grid, $E_{1}$ is the set of ’horizontal’ edges and $E_{2}$ is the set of ’vertical’ edges then $lsimw((G[E_{1}])=lsimw(G[E_{2}])=1$, while $lsimw(G)=\Omega(n)$. Nevertheless, Lemma 8 states that for each linear order $\pi$ of $V(G)$ must be $\Omega(n)$ either for $G[E_{1}]$ or for $G[E_{2}]$. Continuing on the discussion in Remark 4, the above phenomenon is precisely the reason of rigidity of obdd-based models representing cnfs with junction variables. We will see in the proof of Theorems 5 and 6 below how this leads to exponential lower bounds.
Next, we state two theorems, whose proof are provided in Section 3.3, that serve as engines for proving lower bounds for Theorems 4, 5, and 6. We will then use Lemmas 7 and Lemma 8 for ’harnessing’ these engines in order to prove the theorems.
@@
Let $B$ be a $\wedge_{d}$-obdd representing $\psi(G)$. Let $\pi^{*}$ be a linear order of $\mathsf{var}(\psi(G))$ obeyed by $B$. Let $M$ be a matching of $G^{*}$ neatly crossing $\pi^{*}$. Then $|B|\geq 2^{\Omega(|M|)}$.
@@
Let $B$ be an obdd representing $\varphi(G)$. Let $\pi$ be a linear order of $V(G)$ obeyed by $B$. Let $M$ be a matching of $G$ crossing $\pi$. Then $|B|\geq 2^{\Omega(|M|)}$.
@@
(Theorem 4) For a sufficiently large $k$, let ${\bf G}_{k}$ be a class of graphs $G_{k}$ of treewidth at most and lsim-width $\Omega(k\log n)$. Existence of such a class of graphs follows Lemma 1. In order to prove the theorem it is sufficient to demonstrate that $\wedge_{d}$-obdd representing $\psi(G_{k})$ is of size $n^{\Omega(k)}$ and that $itw(\psi(G))=O(k)$.
Let $B$ be a $\wedge_{d}$-obdd representing $\psi(G_{k})$. Let $\pi^{*}$ be a linear order of $V(G^{*}_{k})$ obeyed by $B$. By Lemma 7, there is a matching $M$ of $G^{*}_{k}$ of size $\Omega(lsimw(G_{k}))$ neatly crossing $\pi^{*}$. Hence, by Theorem 9, the size of $2^{\Omega(lsimw(G))}$ which is $n^{\Omega(k)}$ according to the previous paragraph.
For the proof that $itw(\psi(G))=O(k)$, let $(T,{\bf B})$ be a tree decomposition of $G$ of width at most $k$. For each $t\in V(T)$, replace each $v\in{\bf B}(t)$ with $v[1]$ and $v[2]$. It is not hard to see that we obtain a tree decomposition for the primal graph $\varphi(G^{*})$. As the largest bag size of $(T,{\bf B})$ is at most $k+1$, the largest bag size of the new tree decomposition is at most $22k+$ 2 and hence $ptw(\varphi(G^{*}))$ is at most $12k+$ 1. By Proposition 2 $itw(\varphi(G^{*}))$ is at most $12k+$ 1. Take a tree decomposition witnessing $itw(\varphi(G^{*}))$ and add to each bag two vertices corresponding to the long clauses. As a result we obtain a tree decomposition of the incidence graph $\psi(G)$ of width at most $32k+$ 3. ∎
We are now turning to the proof of Theorems 5 and 6. For both theorems, the target class of graphs are $n\times n$ grids that we denote by $G_{n}$. A formal definition is provided below.
@@
Let $n>0$ be an integer. A $n\times n$ grid denoted by $G_{n}$ is a graph with $V(G_{n})=\{(i,j,)|1\leq i\leq n,1\leq j\leq n\}$ and $E(G_{n})=E_{hor}\cup E_{vert}$ where $E_{hor}=\{\{(i,j),(i,j+1)\}|1\leq i\leq n,1\leq j\leq n-1\}$ and $E_{vert}=\{\{(i,j),(i+1,j\}|1\leq i\leq n-1,1\leq j\leq n\}$. We also refer to $E_{hor}$ and $E_{vert}$ as, respectively, horizontal and vertical edges of $G_{n}$.
@@
Let $B$ be a $\wedge_{d}$-obdd obeying an order $\pi$. Let ${\bf g}=\{(x,i)\}$ where $x\in var(B)$ and $i\in\{0,1\}$. Suppose that all the variables of $f(B)|_{\bf g}$ are essential. Let $B^{\prime}$ be obtained from $B$ by the following operations.
@@
    Removal of all the edges $(u,v)$ such that $u$ is a decision node associated with $x$ and $(u,v)$ is labelled with $2-i$.
@@
Then $B^{\prime}$ is a $\wedge_{d}$-obdd obeying $\pi$ and $f(B^{\prime})=f(B)|_{\bf g}$.

### v2

@@ 3.1 An approach to proving $\wedge_{d}$-OBDD lower bounds
Let the notation and premises be as in Lemma 2. In addition, assume that $\mathcal{S}(B)|_{\bf g}\neq\emptyset$. Let $X=(\mathsf{var}(B)\setminus\mathsf{var}({\bf g}))\setminus\bigcup_{u\in L({\bf g})}\mathsf{var}(B_{u})$. (In other words, $X$ is the set of variables not used by the model under the restriction posed by $\bf g$}.) If $L({\bf g})\neq\emptyset$ then $\mathcal{S}(B)|_{\bf g}=X^{\{0,1\}}\times\prod_{u\in L({\bf g})}\mathcal{S}(B_{u})$. Otherwise, $\mathcal{S}(B)|_{\bf g}=X^{\{0,1\}}$.
We believe that Figure 3 makes the statement of Lemma 3 intuitively clear. We formally prove the lemma by induction on $|\mathsf{var}(B)|$. The proof is rather tedious due to the need to inductively account $L({\bf g})$ and $X$ as composed from the corresponding objects for subgraphs of $B$ and the need to prove that the induction assumption holds for these subgraphs. We therefore postpone the full proof to the Appendix and proceed to the use of the lemma for the proposed method of proving lower bounds for $\wedge_{d}$-obdds.
The central part of the method is the notion of a fooling set of a Boolean function.
Definition 10.
Let $f$ be a Boolean function and let $\pi$ be a permutation of $\mathsf{var}(f)$. Let $\mathcal{F}$ be a set of assignments over a prefix $\pi_{0}$ of $\pi$. We say that $\mathcal{F}$ is a ($\wedge_{d}$-obdd-)fooling set for $f$ under $\pi$ if for each $\bf g$$\in\mathcal{F}$ there is an unbreakable set $UB({\bf g})\subseteq\pi\setminus\pi_{0}$ of size at least two having the following properties.
  1. 1.
    Unbreakability: For each $\bf g$$\in\mathcal{F}$, $\mathcal{S}(f)$ does not break $UB({\bf g})$.
  2. 2.
    Distinct projections: For every two distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$, $Proj(\mathcal{S}(f)|_{{\bf g}_{1}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))\neq Proj(\mathcal{S}(f)|_{{\bf g}_{2}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))$.
In the next lemma, we demonstrate that if $\mathcal{F}$ is a fooling set of a Boolean function $f$ under a permutation $\pi$ then any $\wedge_{d}$-obdd representing $f$ and respecting $\pi$ must have size at least $|\mathcal{F}|$. The proof consists of two stages. At the first stage, we consider an arbitrary $\bf g$$\in\mathcal{F}$. Combining the unbreakability property with Lemma 3, we identify precisely one $u\in L({\bf g})$ such that $UB({\bf g})\subseteq\mathsf{var}(B_{u})$. We refer to this $u$ as $u({\bf g})$.
In the second part of the proof, we demonstrate that the mapping from $\bf g$} to $u({\bf g})$ is injective thus implying the desired lower bound. For that purpose, we use the distinct projections property. In particular, for the sake of contradiction, we assume that there are two distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$ such that $u({\bf g}_{1})$ and $u({\bf g}_{2})$ is the same node and denote this node by $u$. Another application of Lemma 3 demonstrates that, for each $i\in[2]$, the projection of $\mathcal{S}(f)|_{{\bf g}_{i}}$ to $UB({\bf g}_{1})\cup UB({\bf g}_{2})$ is $Proj(\mathcal{S}(B_{u}),UB({\bf g}_{1})\cup UB({\bf g}_{2}))$. That is, both projections are the same in contradiction to the distinct projections property.
Now, let us formally state and prove the lemma.
Lemma 4.
Let $f$ be a Boolean function and let $\mathcal{F}$ be a fooling set for $f$ under a permutation $\pi$ of $\mathsf{var}(f)$. Let $B$ be a $\wedge_{d}$-obdd representing $f$ and respecting $\pi$. Then $|B|\geq|\mathcal{F}|$.
Proof.
First, we note that $\mathcal{S}(B)=\mathcal{S}(f)$. Let $\bf g$$\in\mathcal{F}$. By Proposition 1 combined with Lemma 3, either $UB({\bf g})\subseteq X$ or there is $u\in L({\bf g})$ with $UB({\bf g})\subseteq\mathsf{var}(B_{u})$. The former is ruled out simply by contradiction to unbreakability ($X$ can be arbitrarily broken whereas $UB({\bf g})$ is of size at least $2$ so $UB({\bf g})\subseteq X$ would mean that breaking across $UB({\bf g})$ is possible). We denote the $u$ such that $UB({\bf g})\subseteq\mathsf{var}(B_{u})$ by $u({\bf g})$. We prove that the mapping from $\mathcal{F}$ to $V(B)$ such that $\bf g$$\in\mathcal{F}$ is mapped to $u({\bf g})$ is injective and this implies the desired lower bound.
Indeed, assume the opposite. Then there are distinct ${\bf g}_{1},{\bf g}_{2}\in\mathcal{F}$ such that $u=u({\bf g}_{1})=u({\bf g}_{2})$. By Lemma 3, for each $i\in[2]$, $Proj(\mathcal{S}(B)|_{{\bf g}_{i}},\mathsf{var}(B_{u}))=\mathcal{S}(B_{u})$. This in particular, implies that for each $i\in[2]$, $Proj(\mathcal{S}(B)|_{{\bf g}_{i}},UB({\bf g}_{1})\cup UB({\bf g}_{2}))=Proj(\mathcal{S}(B_{u}),UB({\bf g}_{1})\cup UB({\bf g}_{2}))$ in contradiction to the distinct projections assumption. ∎
Thus, in order to prove that a $\wedge_{d}$-obdd representing a function $f$ and respecting a specific order $\pi$ is large, it is enough to demonstrate existence of a large fooling set for $f$ under $\pi$. If $\pi$ is chosen arbitrary, this also implies a lower bound for an arbitrary $\wedge_{d}$-obdd representing $f$. The advantage of this approach is that it is completely ‘function-based’ and does not require direct reasoning about the model. Of course, proving existence of a large $\wedge_{d}$-fooling set may also be challenging. However, the difficulty is comparable with the difficulty of establishing a large fooling set for obdds. Indeed, in the latter case, we effectively need to establish the distinct projection property whereas, in the former case, we also need to establish existence of unbreakable sets. We believe that this extra layer of difficulty is reasonable since $\wedge_{d}$-obdd is a much more powerful model than obdd (as demonstrated in the next subsection). Thus, the proposed approach is a useful tool for establishing lower bounds. We demonstrate its work in Section 3.3 where we use the tool for proving Theorem 10.
@@ 3.2 $\wedge_{d}$-OBDD lower bounds and separations
In [5] we established an xp lower bound for $\wedge_{d}$-obdds representing cnfs of bounded incidence treewidth. We prove this result here using the approach presented in Section 3.1.
@@
[…]
Thus a natural restriction on $\wedge_{d}$-fbdd separates primal [14] and incidence treewidth. This makes $\wedge_{d}$-obdd an interesting object to study in greater detail. We establish the relationship of $\wedge_{d}$-obdd with two closely related formalisms obdd and fbdd by proving exponential separations as specified below (for both statements, the lower bounds are proved using the approach as in Section 3.1). In the next section, we study the Apply operation for $\wedge_{d}$-obdds.
@@
For the corresponding quasipolynomial separation, take a class of cnfs of ptwptw at most $k$ that require fbdd of size $n^{\Omega(k)}$ (see [19] for such a class of cnfs). Let $k=\log n$ where $n$ is the number of variables. Then the resulting class of cnfs will require $n^{\Omega(\log(n)}$ size fbdds for their representation but can be represented by a polynomial size $\wedge_{d}$-obdd according to [14].
@@
Theorem 7 looks somewhat surprising because equipping fbdds with the $\wedge_{d}$-nodes leads only to quasipolynomial gains in size [3]
@@
$E(G)=\{\{u_{1},u_{2}\},\{u_{2},u_{3}\},\{u_{3},u_{4}\},\{u_{4},u_{5}\},\{u_{5},u_{1}\},\{u_{1},u_{3}\}\}$. In other words, $G$ is a cycle of $5$ vertices with an extra chord. Then $V(G^{*})=\{u_{1}[1],\dots,u_{5}[1],u_{1}[2],\dots,u_{5}[2]\}$ and $E(G^{*})=\{\{u_{1}[1],u_{2}[2]\},\{u_{2}[1],u_{3}[2]\},\{u_{3}[1],u_{4}[2]\},\{u_{4}[1],u_{5}[2]\},\{u_{5}[1],u_{1}[2]\},\{u_{1}[1],u_{3}[2]\},\\ \{u_{1}[2],u_{2}[1]\},\{u_{2}[2],u_{3}[1]\},\{u_{3}[2],u_{4}[1]\},\{u_{4}[2],u_{5}[1]\},\{u_{5}[2],u_{1}[1]\},\{u_{1}[2],u_{3}[1]\}\}$. $\varphi(G)$ and $\varphi(G^{*})$ simply turn the edges $\{u,v\}$ of the respective graphs into clauses $(u\vee v)$. $\psi(G)$ is obtained from $\varphi(G^{*})$ by adding clauses $(\neg u_{1}[1]\vee\dots\vee\neg u_{5}[1])$ and $(\neg u_{1}[2]\vee\dots\vee\neg u_{5}[2])$
@@
We will use cnfs with junction variables for Theorems 6 and 7. This approach is used in [25] for separation of obdds and fbdds. The rough idea is that a ’hard’ function is ’split’ into two ’easy’ ones using the junction variable. The fbdd, due to its flexibility, can adjust and represent each ’easy’ function with the suitable variable ordering. On the other hand, obdd is rigid in the sense that it needs to use the same ordering for both ’easy’ functions and the ordering leads to an exponential size representation for one of them.
In order to proceed, we need to delve deeper into the structure of matchings of the graph $G^{*}$. This is done in the next definition and the two statements that follow.
@@
Let $M$ be a matching of $G^{*}$ between $V^{\prime}$ and $V^{\prime\prime}$. We say that $V^{\prime},V^{\prime\prime}$ is the neat partition of $V(M)$ if there are $U,W\subseteq V(G)$ such that $V^{\prime}=U[1]$ and $V^{\prime\prime}=W[2]$.
Let $\pi^{*}$ be a linear order of $V(G^{*})$. We say that $M$ neatly crosses $\pi^{*}$ if there a witnessing prefix $\pi^{*}_{0}$ of $\pi^{*}$ such that $\pi^{*}_{0}\cap V(M)$ and $(\pi^{*}\setminus\pi^{*}_{0})\cap V(M)$ form a neat partition of $M$.
Example 5.
Consider a graph $G$ with four vertices $u^{\prime},u^{\prime\prime},v^{\prime},v^{\prime\prime}$ and two edges: $\{u^{\prime},u^{\prime\prime}\}$ and $\{v^{\prime},v^{\prime\prime}\}$. Let $M=\{\{u^{\prime}[1],u^{\prime\prime}[2]\},\{v^{\prime\prime}[1],v^{\prime}[2]\}\}$. Clearly, $M$ is a matching of $G^{*}$. The neat partition for $M$ is $\{\{u^{\prime}[1],v^{\prime\prime}[1]\},\{u^{\prime\prime}[2],v^{\prime}[2]\}\}$, the corresponding partition of $V(G)$ is $\{\{u^{\prime},v^{\prime\prime}\},\{u^{\prime\prime},v^{\prime}\}\}$.
Consider the linear order $\pi^{*}=(v^{\prime\prime}[2],u^{\prime}[1],u^{\prime}[2],u^{\prime\prime}[1],v^{\prime\prime}[1],v^{\prime}[1],u^{\prime\prime}[2],v^{\prime}[2])$. The order neatly crosses $M$ as witnessed by the prefix $\pi^{*}_{0}=(v^{\prime\prime}[2],u^{\prime}[1],u^{\prime}[2],u^{\prime\prime}[1],v^{\prime\prime}[1])$. On the other hand, if in $\pi^{*}$ we swap $v^{\prime\prime}[1]$ and $u^{\prime\prime}[2]$ then we obtain a linear order of $V(G^{*})$ that does not cross $M$ neatly.
In the next statement, we demonstrate that if each linear order of $G$ is crossed by a large induced matching then every linear order of $G^{*}$ is neatly crossed by a large induced matching. The statement is formulated in terms of parameter lsim-width defined earlier.
@@
Every linear order of $G^{*}$ is neatly crossed by an induced matching of size at least $lsimw(G)/2$
@@
Let $u[1]$ be the last vertex of $\pi^{*}_{0}$. Let $w[2]\in W_{1}[2]$. Note that by definition of $\pi_{0}$, $u\in\pi_{0}$ while $w$ is not. It follows from $u[ind(u)]<w[ind(w)]$ according to $\pi^{*}$. Now $ind(u)=1$ by definition of $U_{1}$. That is $u[1]$ occurs before $w[ind(w)]$ and surely before w[3ind(w)]w[3-ind(w)] in $\pi^{*}$. As $u[1]$ is the last element of $\pi^{*}_{0}$, both occurrences of $w$ are outside of $\pi^{*}_{0}$. ∎
@@
For $i\in\{1,2\}$, Let $E^{*}_{i}=E(G[E_{i}]^{*})$. It is not hard to see that $E^{*}_{1},E^{*}_{2}$ partition $E(G^{*})$ and $G^{*}[E^{*}_{i}]=G[E_{i}]^{*}$ for each $i\in\{1,2\}$. Let $M^{*}$ be an induced matching of size at least $lsimw(G)/2$ neatly crossing $\pi^{*}$ existing by Lemma 8. For each $i\in\{1,2\}$, let $M^{*}_{i}=M^{*}\cap E^{*}_{i}$. Clearly, both $M^{*}_{1}$ and $M^{*}_{2}$ cross $\pi^{*}$ and least one of them is of size at least half the size of $M^{*}$ that is at least $lsimw(G)/4$. ∎
@@
Note that Lemma 9 does not make any statements regarding $lsimw(G[E_{1}])$ and $lsimw(G[E_{2}])$. Indeed, if $G$ is an $n\times n$ grid, $E_{1}$ is the set of ’horizontal’ edges and $E_{2}$ is the set of ’vertical’ edges then $lsimw((G[E_{1}])=lsimw(G[E_{2}])=1$, while $lsimw(G)=\Omega(n)$. Nevertheless, Lemma 9 states that for each linear order $\pi$ of $V(G)$ must be $\Omega(n)$ either for $G[E_{1}]$ or for $G[E_{2}]$. Continuing on the discussion in Remark 5, the above phenomenon is precisely the reason of rigidity of obdd-based models representing cnfs with junction variables. We will see in the proof of Theorems 6 and 7 below how this leads to exponential lower bounds.
Next, we state two theorems, whose proofs are provided in Section 3.3, that serve as engines for proving lower bounds for Theorems 5, 6, and 7. We will then use Lemmas 8 and Lemma 9 for ‘harnessing’ these engines in order to prove the theorems. For both of the theorems below, $G$ is a graph without isolated vertices.
@@
Let $B$ be a $\wedge_{d}$-obdd representing $\psi(G)$. Let $\pi^{*}$ be a linear order of $\mathsf{var}(\psi(G))$ respected by $B$. Let $M$ be an induced matching of $G^{*}$ neatly crossing $\pi^{*}$. Then $|B|\geq 2^{\Omega(|M|)}$.
@@
Let $B$ be an obdd representing $\varphi(G)$. Let $\pi$ be a linear order of $V(G)$ respected by $B$. Let $M$ be an induced matching of $G$ crossing $\pi$. Then $|B|\geq 2^{\Omega(|M|)}$.
@@
(Theorem 5) For a sufficiently large $k$, let ${\bf G}_{k}$ be a class of graphs $G_{k}$ of treewidth at most $k$ and lsim-width $\Omega(k\log n)$. Existence of such a class of graphs follows from Lemma 1. In order to prove the theorem it is sufficient to demonstrate that $\wedge_{d}$-obdd representing $\psi(G_{k})$ is of size $n^{\Omega(k)}$ and that $itw(\psi(G))=O(k)$.
Let $B$ be a $\wedge_{d}$-obdd representing $\psi(G_{k})$. Let $\pi^{*}$ be a linear order of $V(G^{*}_{k})$ obeyed by $B$. By Lemma 8, there is an induced matching $M$ of $G^{*}_{k}$ of size $\Omega(lsimw(G_{k}))$ neatly crossing $\pi^{*}$. Hence, by Theorem 10, the size of $B$ is $2^{\Omega(lsimw(G))}$ which is $n^{\Omega(k)}$ according to the previous paragraph.
For the proof that $itw(\psi(G))=O(k)$, let $(T,{\bf B})$ be a tree decomposition of $G$ of width at most $k$. For each $t\in V(T)$, replace each $v\in{\bf B}(t)$ with $v[1]$ and $v[2]$. It is not hard to see that we obtain a tree decomposition for the primal graph $\varphi(G^{*})$. As the largest bag size of $(T,{\bf B})$ is at most $k+1$, the largest bag size of the new tree decomposition is at most $2k+2$ and hence $ptw(\varphi(G^{*}))$ is at most $2k+1$. By Proposition 2 $itw(\varphi(G^{*}))$ is at most $2k+1$. Take a tree decomposition witnessing $itw(\varphi(G^{*}))$ and add to each bag two vertices corresponding to the long clauses. As a result we obtain a tree decomposition of the incidence graph $\psi(G)$ of width at most $2k+3$. ∎
We are now turning to the proof of Theorems 6 and 7. For both theorems, the target class of graphs are $n\times n$ grids that we denote by $G_{n}$. A formal definition is provided below.
@@
Let $n>0$ be an integer. A $n\times n$ grid, denoted by $G_{n}$, is a graph with $V(G_{n})=\{(i,j)|1\leq i\leq n,1\leq j\leq n\}$ and $E(G_{n})=E_{hor}\cup E_{vert}$ where $E_{hor}=\{\{(i,j),(i,j+1)\}|1\leq i\leq n,1\leq j\leq n-1\}$ and $E_{vert}=\{\{(i,j),(i+1,j)\}|1\leq i\leq n-1,1\leq j\leq n\}$. We also refer to $E_{hor}$ and $E_{vert}$ as, respectively, horizontal and vertical edges of $G_{n}$.
@@
Let $B$ be a $\wedge_{d}$-obdd respecting an order $\pi$. Let ${\bf g}=\{(x,i)\}$ where $x\in\mathsf{var}(B)$ and $i\in\{0,1\}$. Suppose that all the variables of $f(B)|_{\bf g}$ are essential.  [^1] A variable $x$ of a Boolean function $h$ is essential is there are two assignments ${\bf a}_{1}$ and ${\bf a}_{2}$ that differ only in their assignment of $x$ such that $h({\bf a}_{1})\neq h({\bf a}_{2})$. Let $B^{\prime}$ be obtained from $B$ by the following operations.
@@
    Removal of all the edges $(u,v)$ such that $u$ is a decision node associated with $x$ and $(u,v)$ is labelled with 1i1-i.
@@
Then $B^{\prime}$ is a $\wedge_{d}$-obdd respecting $\pi$ and $f(B^{\prime})=f(B)|_{\bf g}$.